OpenStudy (anonymous):

A toy manufacturer needs a piece of plastic in the shape of a right triangle with the longer leg 1 cm more than the shorter leg and the hypotenuse 2 cm more than the shorter leg. How long should the sides of the triangle be?

OpenStudy (anonymous):

 pythagorean theorem

OpenStudy (mathmate):

You probably guessed the answer. If you want to proceed with a formal solution, it would be: Solve for x in x^2+(x+1)^2=(x+2)^2 Expand both sides and regroup terms: x^2-2x-3=0 (x-3)(x+1)=0 So there...

OpenStudy (anonymous):

3,-1??